# An independent search for Jovian neutrinos using BOREXINO data

### v1

@@ Abstract
In a recent study,  Ansarifard and Farzan (2024) found evidence for a 6% flux contribution from Jupiter to the total flux rate time series data from the BOREXINO solar neutrino experiment, specifically during the time intervals 2019-2021 and 2011-2013. The significance of this detection was estimated to be around $2\sigma$. We reanalyze the BOREXINO data and independently confirm the Jovian signal with the same amplitude and significance as that obtained in  Ansarifard and Farzan (2024). However, using the same technique, we also find a spurious flux contribution from Venus and Saturn (at $\sim 2\sigma$ significance), whereas prima facie one should not expect any signal from any other planet. We then implement Bayesian model comparison to ascertain whether the BOREXINO data contain an additional contribution from Jupiter, Venus or Saturn. We find Bayes factors of less than five for an additional contribution from Jupiter, and less than or close to one for Venus and Saturn. This implies that the evidence for an additional contribution from Jupiter is very marginal.
@@ I Introduction
In a recent work,  Ansarifard and Farzan (2024) (A24 hereafter) found that the temporal variation in the 7Be solar neutrino signal from the BOREXINO experiment exhibits a modulation, which is attributed to a contribution from Jupiter at about $2\sigma$ significance. The flux from Jupiter has been estimated to be about 6% of the total solar neutrino signal observed. This Jovian contribution was also able to explain the suppression in the time variation of the signal as well as the lower value of the extracted eccentricity. It was subsequently argued that this signal is caused by capture and annihilation of dark matter WIMPs of masses less than 4 GeV in Jupiter. If this signal is independently confirmed, this would be the first evidence of neutrinos of MeV energy detected from any other solar system object other than the Sun. Although, a large number of detectors, starting with the Homestake experiment in 1960s Davis Jr et al. (1968) have measured the solar neutrino flux, no other experiment has found any signature for Jovian neutrinos. In a similar vein, although neutrino detectors have been searching for signatures of dark matter capture and annihilation for a long time Desai et al. (2004), this would possibly be the first signature of WIMP annihilation to neutrinos from any solar system object.
Therefore, given the profound implications of this result, it behooves us to carry out an independent search for Jovian neutrinos in the BOREXINO data. We therefore follow the same prescription as in A24 and supplement the analysis using Bayesian model comparison. We then also do a similar search from some other planets in order to compare the results with those from Jupiter, as a null test of the analysis method used.
This manuscript is structured as follows. We describe the BOREXINO data used for this analysis in Sect. II. Our results using the total rate time series data and modulation data are presented in Sect. III and Sect. IV, respectively. We conclude in Sect. V.
@@ II Summary of A24
where $\mathcal{R}_{sun}$ and $\mathcal{R}_{jup}$ are proportional to the event rates induced by Sun and Jupiter, respectively, and $\mathcal{R}_{B}$ is the background contribution from radioactive sources. In Eq 1, $d_{\mathrm{sun}}$ and $d_{\mathrm{jup}}$ denote the instantaneous distance to the Sun and Jupiter, respectively. A24 assumed that $\mathcal{R}_{sun}$ and $\mathcal{R}_{jup}$ are constant. A24 then fit the data between October 2019 and October 2021 to Eq 1 using Bayesian regression, when the temporal variation of $\mathcal{R}_{B}$ was negligible Borexino Collaboration et al. (2020). For Bayesian inference, uniform priors for $\mathcal{R}_{jup}$ and $\mathcal{R}_{B}$ were assumed, both of which given by $\in\mathcal{U}[0,50]$. For $\mathcal{R}_{sun}$ two sets of priors were used: $\mathcal{U}[0,50]$ and $\mathcal{N}(25,2)$. A24 obtained the best-fit value of $\frac{\mathcal{R}_{jup}}{(5\ {\rm AU})^{2}}=1.5^{+0.7}_{-0.8}\ (\rm cpd/100t)$, which points to non-zero value at $2\sigma$.
@@
where the index $i$ represents the two objects considered, Sun and Jupiter and $T$ is the duration of the dataset analyzed.  [^2] Equation (6) in A24 does not contain $T$, which is a typo (S. Ansarifard, private communication). A24 found no contribution from Jupiter during 2015-2018. However, non-zero values were found during two distinct periods: December 2011-December 2013 and October 2019-October 2021 with best-fit values given by $\frac{\mathcal{R}_{jup}}{(5\ {\rm AU})^{2}}=1.6^{+0.8}_{-1.1}\ (\rm cpd/100t)$ and $\frac{\mathcal{R}_{jup}}{(5\ {\rm AU})^{2}}=1.7^{+0.8}_{-0.8}\ (\rm cpd/100t)$, respectively. Therefore, both method yield a signal which is about 6% of the total 7Be solar neutrino flux obsered.
Finally, A24 also repeated the same exercise on data from SNO annual variation data from 1999-2003 in the energy range from 5-20 MeV Aharmim et al. (2005), but no signal was found.
@@ III.1 Results of Bayesian model comparison
[…]
@@
For our analysis, the model $M_{2}$ corresponds to the hypothesis that the BOREXINO rate time-series data are a superposition of flux from Jupiter and Sun, in addition to the background from radioactivity, while $M_{1}$ corresponds to the hypothesis that the rate time-series data is due to a combination of solar neutrino flux and radioactive background. For this analysis, similar to Bayesian inference, we use a Gaussian likelihood and the same priors as A24. To calculate the Bayesian evdience and Bayes factors, we use the Dynesty nested sampler Speagle (2020). The results for these Bayes factors can be found in Table 2. We find for all ephemerides, that for a Gaussian prior on $\mathcal{R}_{sun}$, the Bayes factor is $\sim 3$ and $\sim 1$ for a uniform prior on $\mathcal{R}_{sun}$. According to Jeffreys’ scale, this value corresponds to “barely worth mentioning” Trotta (2017), implying that the evidence for the extra contribution from Jupiter is marginal. Therefore, Bayesian model comparison provides inconclusive evidence for a contribution from Jupiter in the BOREXINO time series data, compared to a model with no such contribution.
@@ III.2 Search for a signal from other planets
We now carry out a similar exercise as in the previous section for other nearby solar system planets, in order to compare and contrast our results with Jupiter. This would also serve as a null test of the Bayesian regression analysis technique used in A24, since apriori we do not expect any signal from any other planets if an additional signal is argued to come from only Jupiter. For this purpose, we repeat the Bayesian regression and model comparison analysis for Venus, Mars and Saturn using the time period between October 2019 and October 2021. We did not consider Mercury since it is very close to the Sun. For Bayesian regression, we replace the term for Jupiter in Eq 1 with the corresponding term for a given planet. Since the average distance to Venus, Mars, and Saturn between 2019-2021 is equal to 1.0, 2.0, and 10.0 AU, respectively, we normalize the signal contributions by the square of these distances similar to that done for Jupiter. We use the same priors for $\mathcal{R}_{\mathrm{sun}}$ as those used for Jupiter. However, we changed the lower limit on $\mathcal{R}_{\mathrm{mar}}$ (Mars), $\mathcal{R}_{\mathrm{ven}}$ (Venus), and $\mathcal{R}_{\mathrm{sat}}$ (Saturn) to values less than zero, in order to see if the results are consistent with null results. We have considered only one ephemeris, namely DE442s. We have also changed the lower limits on the prior for each of the planets to extend to negative values, in order to check if the signal is compatible with zero value.
@@
The corresponding marginalized 68% and 90% credible intervals can be found in Figs. 3 and  4, Figs. 5 and  6, and Figs. 7 and  8 respectively for Venus, Mars, and Saturn respectively. A tabular summary can be found in Table 3. For Mars, the $1\sigma$ central intervals are consistent with zero. For Saturn, we get $1.6\sigma$ evidence for non-zero flux for a Gaussian prior on $\mathcal{R}_{\mathrm{sun}}$ with values of $\frac{\mathcal{R}_{\mathrm{sat}}}{(10\,\mathrm{AU})^{2}}=2.59^{+1.59}_{-1.62}$ (cpd/100t). However, for a uniform prior on $\mathcal{R}_{\mathrm{sun}}$, the contribution from Saturn is consistent with zero. For Venus, we get central estimates for $\frac{\mathcal{R}_{\mathrm{ven}}}{(1\,\mathrm{AU})^{2}}$ equal to $0.18^{+0.09}_{-0.09}$ and $0.16^{+0.09}_{-0.09}$ (cpd/100t), for Gaussian and uniform priors on $\mathcal{R}_{\mathrm{sun}}$, corresponding to about $2\sigma$ significances, similar to that seen for Jupiter. The inferred flux from Venus about 10% that of Jupiter and about 0.6% compared to the solar contribution. Therefore, the Bayesian regression technique implemented in A24 also returns spurious signals for Venus ($2\sigma$) and Saturn ($1.6\sigma$) (for one of the prior choices).
@@
We also carried out Bayesian model comparison for these planets similar to that done for Jupiter. Similar to before, the null hypothesis involves the BOREXINO rate time-series data containing a contribution from only Sun in addition to the radioactive backgrounds, whereas the alternative hypothesis involves an additional contribution from the corresponding planet. Unlike Bayesian regression, we used uniform priors with only a positive contribution from a given planet, with the minimum value of 0, since we only want to compare the significance of a positive signal. The results for the Bayes factors can be found in Table 4, along with the priors used. We find that the Bayes factors are less than 1 for Venus and Mars with values $\lessapprox 0.01$, implying that additional contributions from Venus and Mars are decisively ruled out. For Saturn we find that the Bayes factors are close to 1, implying that both the significance of both the hypotheses are comparable.
@@
Therefore, Bayesian regression also provides a $2\sigma$ evidence for a (spurious) contribution from Venus and $1.6\sigma$ contribution from Saturn (for a Gaussian prior on Sun) similar to that from Jupiter. However a Bayesian model comparison does not support any additional contribution to the total rate from Mars, Venus, and Saturn.
@@
Mars $\frac{\mathcal{R}_{\mathrm{mar}}}{(2\,\mathrm{AU})^{2}}$ $\mathcal{U}[-13,13]$ $0.02^{+0.04}_{-0.04}$ $0.02^{+0.04}_{-0.04}$
@@ III.3 Goodness of fit tests
Therefore, we find that Bayesian regression points to evidence for additional flux from Jupiter, Venus, and Saturn (for one chosen prior), whereas, Bayesian model comparison does not support any additional contribution from Jupiter, Venus or Saturn. We then overlay the total time series data to all three models considered so far, involving a contribution from only Sun, Sun+Jupiter, Sun+Venus, and Sun+Saturn This plot can be found in Fig. 9. We also calculate the reduced $\chi^{2}$ for all four hypotheses, to check if each of these provides a good fit. We find $\chi^{2}/dof$ equal to 28.6/22 (1.3), 24.3/21 (1.15), 25.3/21 (1.2), 26.2/21 (1.2) for Sun, Sun+Venus, Sun+Jupiter, Sun+Saturn respectively. Therefore, the reduced $\chi^{2}$ are close to one for all the four hypotheses.
@@ IV Analysis with monthly modulation data
$\delta\mathcal{R}(t)=\sum_{i}\mathcal{R}_{i}\left[\frac{1}{d^{2}_{i}(t)}-\biggl{<}\frac{dt}{d^{2}_{i}(t)}\biggr{>}\right],$ (5)
where the index $i$ once again runs over Sun and Jupiter. The best-fit marginalized values for $\mathcal{R}_{\mathrm{jup}}$ for the two time intervals can be found in Fig. 10 for the DE422s ephemerides. These are in agreement with Fig. 3 of A24. The best-fit values for $\mathcal{R}_{\mathrm{jup}}$ are equal to $1.78^{+1.03}_{-0.97}$ and $1.69^{+0.79}_{-0.77}$ (cpd/100t), for 2011-2013 and 2019-2021 respectively, corresponding to significances of $1.8\sigma$ and $2.1\sigma$. We also did a similar exercise using other ephemerides. These values can be found in Table 5. We find that results for both epochs are consistent with DE422s for all the ephemerides used. Therefore, even with the monthly modulation data we find evidence for $2\sigma$ contribution from Jupiter with flux contribution of about 6% of that seen from the Sun.
We then do Bayesian model comparison in the same way as done in Sect. III. We used a Gaussian likelihood and the same prior for $\mathcal{R}_{\mathrm{sun}}$ and $\mathcal{R}_{\mathrm{jup}}$ as used in the Bayesian regression. The null hypothesis corresponds to the ansatz that the modulation data is due to only the Sun, whereas the alternate hypothesis is that the modulation data is a combination of data from Sun and Jupiter. The Bayes factors for all the ephemerides for both epochs can be found in Table 6. The Bayes factor for 2011-2013 period is around two and hence not significant. Although the Bayes factor for the 2019-2021 period are marginally larger than with the total rate time-series data with value of about four, they are still $<5$ and according to Jeffreys scale, still point to “barely worth mentioning”. Therefore, a Bayesian model comparison still only provides a marginal support for an additional contribution from Jupiter to the monthly modulation data during the periods 2019-2021 and 2011-2013.

### v2

@@ Abstract
In a recent study,  Ansarifard and Farzan (2024) found evidence for a 6% flux contribution from Jupiter to the total rate time series data from the BOREXINO solar neutrino experiment, specifically during the time intervals 2019-2021 and 2011-2013. The significance of this detection was estimated to be around $2\sigma$. We reanalyze the BOREXINO data and independently confirm the Jovian signal with the same amplitude and significance as that obtained in  Ansarifard and Farzan (2024). However, using the same regression technique, we also find a spurious flux contribution from Venus and Saturn at $\sim 2\sigma$ significance, whereas prima facie one should not expect any signal from any other planet. We then implement Bayesian model comparison to ascertain whether the BOREXINO data contain an additional contribution from Jupiter, Venus, or Saturn. We find Bayes factors of less than five for an additional contribution from Jupiter, and less than or close to one for Venus and Saturn. This implies that the evidence for an additional contribution from Jupiter is very marginal.
@@ I Introduction
In a recent work,  Ansarifard and Farzan (2024) (A24 hereafter) found that the temporal variation in the 7Be solar neutrino signal from the BOREXINO experiment exhibits a modulation, which they attributed to a contribution from Jupiter at about $2\sigma$ significance. The flux from Jupiter has been estimated to be about 6% of the total observed solar neutrino signal. This Jovian contribution was also able to explain the suppression in the time variation of the signal as well as the lower value of the Earth’s eccentricity. It was subsequently argued in A24 that this observed signal from Jupiter is caused by capture and annihilation of dark matter WIMPs of masses less than 4 GeV in Jupiter’s core. If this signal is independently confirmed, this would be the first evidence of neutrinos of MeV energy detected from any other solar system object other than the Sun. Although, a large number of detectors, starting with the Homestake experiment in 1960s Davis Jr et al. (1968) have measured the solar neutrino flux, no other detector has found any signature for Jovian neutrinos. In a similar vein, although neutrino detectors have been searching for signatures of dark matter capture and annihilation for a long time Desai et al. (2004), this would possibly be the first signature of WIMP annihilation to neutrinos from any solar system object.
Therefore, given the profound implications of this result, it behooves us to carry out an independent search for Jovian neutrinos in the BOREXINO data. We therefore follow the same prescription as in A24 and supplement the analysis using Bayesian model comparison. We then also do a similar search from some other planets in order to compare the results with those from Jupiter, as a null test of the analysis method used. One should not expect neutrinos from any other planet. If the same analysis technique shows signals from other planets with similar significance as that observed from Jupiter, that would invalidate the results from Jupiter.
This manuscript is structured as follows. We describe the BOREXINO data used for this analysis in Sect. II. Our results using the total rate time series data and modulation data are presented in Sect. III and Sect. IV, respectively. The Jovian analysis using the full data can be found in Sect. V. The results from a frequentist analysis can be found in Sect. VI. We conclude in Sect. VII.
@@ II Summary of A24
where $\mathcal{R}_{sun}$ and $\mathcal{R}_{jup}$ are proportional to the event rates induced by Sun and Jupiter, respectively, and $\mathcal{R}_{B}$ is the background contribution from radioactive sources. In Eq 1, $d_{\mathrm{sun}}$ and $d_{\mathrm{jup}}$ denote the instantaneous distance to the Sun and Jupiter, respectively. A24 assumed that $\mathcal{R}_{sun}$ and $\mathcal{R}_{jup}$ are constant. A24 then fit the data between October 2019 and October 2021 to Eq 1 using Bayesian regression, when the temporal variation of $\mathcal{R}_{B}$ was negligible Borexino Collaboration et al. (2020). For Bayesian inference, uniform priors for $\mathcal{R}_{jup}$ and $\mathcal{R}_{B}$ were assumed, since there is no prior theoretical model for a sigmal from Jupiter. Both these priors are given by $\in\mathcal{U}[0,50]$. For $\mathcal{R}_{sun}$ two sets of priors were used: $\mathcal{U}[0,50]$ and $\mathcal{N}(25,2)$. A24 obtained the best-fit value of $\frac{\mathcal{R}_{jup}}{(5\ {\rm AU})^{2}}=1.5^{+0.7}_{-0.8}\ (\rm cpd/100t)$, which points to non-zero value at $2\sigma$.
@@
where the index $i$ represents the two objects considered, Sun and Jupiter and $T$ is the duration of the dataset analyzed.  [^2] Equation (6) in A24 does not contain $T$, which is a typographical error (S. Ansarifard, private communication). A24 found no contribution from Jupiter during 2015-2018. However, non-zero values were found during two distinct periods: December 2011-December 2013 and October 2019-October 2021 with best-fit values given by $\frac{\mathcal{R}_{jup}}{(5\ {\rm AU})^{2}}=1.6^{+0.8}_{-1.1}\ (\rm cpd/100t)$ and $\frac{\mathcal{R}_{jup}}{(5\ {\rm AU})^{2}}=1.7^{+0.8}_{-0.8}\ (\rm cpd/100t)$, respectively. Therefore, both methods yield a signal which is about 6% of the total 7Be solar neutrino flux obsered.
Finally, A24 also repeated the same exercise on data from Sudbury Neutrino observatory annual variation data from 1999-2003 in the energy range from 5-20 MeV Aharmim et al. (2005), but no signal was found.
@@ III.1 Results of Bayesian model comparison
$B_{21}=\frac{\int P(D|M_{2},\theta_{2})P(\theta_{2}|M_{2})\,d\theta_{2}}{\int P(D|M_{1},\theta_{1})P(\theta_{1}|M_{1})\,d\theta_{1}},$ (4)
@@
For our analysis, the model $M_{2}$ corresponds to the hypothesis that the BOREXINO rate time-series data are a superposition of flux from Jupiter and Sun, in addition to the background from radioactivity, while $M_{1}$ corresponds to the hypothesis that the rate time-series data is due to a combination of only solar neutrino flux and radioactive background. For this analysis, similar to Bayesian inference, we use a Gaussian likelihood and the same priors as A24. To calculate the Bayesian evdience and Bayes factors, we use the Dynesty nested sampler Speagle (2020). The results for these Bayes factors can be found in Table 2. We find that for all ephemerides, the Bayes factor is $\sim 3$ for a Gaussian prior on $\mathcal{R}_{sun}$, and $\sim 1$ for a uniform prior on $\mathcal{R}_{sun}$. According to Jeffreys’ scale, this value corresponds to “barely worth mentioning” Trotta (2017), implying that the evidence for the extra contribution from Jupiter is marginal. Therefore, Bayesian model comparison provides inconclusive evidence for a contribution from Jupiter in the BOREXINO time series data, compared to a model with no such contribution.
@@ III.2 Search for a signal from other planets
We now carry out a similar exercise as in the previous section for other nearby solar system planets, in order to compare and contrast our results with Jupiter. This would also serve as a null test of the Bayesian regression analysis technique used in A24, since apriori we do not expect any signal from any other planets, if an additional signal is argued to come from only Jupiter. For this purpose, we repeat the Bayesian regression and model comparison analysis for Venus, Mars, and Saturn using the time period between October 2019 and October 2021. We did not consider Mercury since it is very close to the Sun and any signal would be degenerate compared to the Sun. For Bayesian regression, we replaced the term for Jupiter in Eq. 1 with the corresponding term for a given planet. Since the average distance to Venus, Mars, and Saturn between 2019-2021 is equal to 1.0, 2.0, and 10.0 AU, respectively, we normalize the signal contributions by the square of these distances similar to that done for Jupiter. We use the same priors for $\mathcal{R}_{\mathrm{sun}}$ as those used for Jupiter. However, we changed the lower limit on $\mathcal{R}_{\mathrm{mar}}$ (Mars), $\mathcal{R}_{\mathrm{ven}}$ (Venus), and $\mathcal{R}_{\mathrm{sat}}$ (Saturn) to values less than zero, in order to see if the results are consistent with zero flux. We have considered only one ephemeris, namely DE442s. We have also changed the lower limits on the prior for each of the planets to extend to negative values, in order to check if the signal is compatible with zero value.
@@
The corresponding marginalized 68% and 90% credible intervals can be found in Figs. 3 and  4, Figs. 5 and  6, and Figs. 7 and  8 respectively for Venus, Mars, and Saturn respectively. A tabular summary can be found in Table 3. For Mars, the $1\sigma$ central intervals are consistent with zero. For Saturn, we get $1.6\sigma$ evidence for non-zero flux for a Gaussian prior on $\mathcal{R}_{\mathrm{sun}}$ with values of $\frac{\mathcal{R}_{\mathrm{sat}}}{(10\,\mathrm{AU})^{2}}=2.59^{+1.59}_{-1.62}$ (cpd/100t). However, for a uniform prior on $\mathcal{R}_{\mathrm{sun}}$, the contribution from Saturn is consistent with zero. For Venus, we get central estimates for $\frac{\mathcal{R}_{\mathrm{ven}}}{(1\,\mathrm{AU})^{2}}$ equal to $0.18^{+0.09}_{-0.09}$ and $0.16^{+0.09}_{-0.09}$ (cpd/100t), for Gaussian and uniform priors on $\mathcal{R}_{\mathrm{sun}}$, corresponding to about $2\sigma$ significances, similar to that seen for Jupiter. The inferred flux from Venus about 10% smaller than Jupiter and about 0.6% compared to the solar contribution. Therefore, the Bayesian regression technique implemented in A24 also returns spurious signals for Venus ($2\sigma$) and Saturn ($1.6\sigma$) (for one of the prior choices).
@@
We also carried out Bayesian model comparison for these planets similar to that done for Jupiter. Similar to before, the null hypothesis involves the BOREXINO rate time-series data containing a contribution from only Sun in addition to the radioactive backgrounds, whereas the alternative hypothesis involves an additional contribution from the corresponding planet. Unlike Bayesian regression, we used uniform priors with only a positive contribution from a given planet, with the minimum value of 0, since we only want to compare the significance of a positive signal. The results for the Bayes factors can be found in Table 4, along with the priors used. We find that the Bayes factors are less than one for Venus and Mars with values $\lessapprox 0.01$, implying that additional contributions from Venus and Mars are decisively ruled out. For Saturn we find that the Bayes factors are close to 1, implying that the significance of both the hypotheses are comparable and there is no support for additional contribution from Saturn.
@@
Therefore, Bayesian regression also provides a $2\sigma$ evidence for a (spurious) contribution from Venus and $1.6\sigma$ contribution from Saturn (for a Gaussian prior on Sun) similar to that from Jupiter. However, a Bayesian model comparison does not support any additional contribution to the total rate from Mars, Venus, and Saturn.
@@
Mars $\frac{\mathcal{R}_{\mathrm{mar}}}{(2\,\mathrm{AU})^{2}}$ $\mathcal{U}[-13,13]$ $02^{+0.04}_{-0.04} 0.0$$02^{+0.04}_{-0.04}$
@@ III.3 Goodness of fit tests
Therefore, we find that Bayesian regression points to $\sim 2\sigma$ evidence for additional flux from Jupiter, Venus, and Saturn (for one chosen prior), whereas, Bayesian model comparison does not support any additional contribution from Jupiter, Venus or Saturn. We then overlay the total time series data to all three models considered so far, involving a contribution from only Sun, Sun+Jupiter, Sun+Venus, and Sun+Saturn This plot can be found in Fig. 9. We also calculate the reduced $\chi^{2}$ for all four hypotheses, to check if each of these provides a good fit. We find $\chi^{2}/dof$ equal to 28.6/22 (1.3), 24.3/21 (1.15), 25.3/21 (1.2), 26.2/21 (1.2) for Sun, Sun+Venus, Sun+Jupiter, Sun+Saturn respectively. Therefore, the reduced $\chi^{2}$ are close to one for all the four hypotheses.
@@ IV Analysis with monthly modulation data
[…]
where the index $i$ once again runs over Sun and Jupiter. The best-fit marginalized values for $\mathcal{R}_{\mathrm{jup}}$ for the two time intervals can be found in Fig. 10 for the DE422s ephemerides. These are in agreement with Fig. 3 of A24. The best-fit values for $\mathcal{R}_{\mathrm{jup}}$ are equal to $1.78^{+1.03}_{-0.97}$ and $1.69^{+0.79}_{-0.77}$ (cpd/100t), for 2011-2013 and 2019-2021 respectively, corresponding to significances of $1.8\sigma$ and $2.1\sigma$. We also did a similar exercise using other ephemerides. These values can be found in Table 5. We find that the results for both epochs are consistent with DE422s for all the ephemerides used. Therefore, even with the monthly modulation data we find evidence for $2\sigma$ contribution from Jupiter with flux contribution of about 6% of that seen from the Sun.
We then do Bayesian model comparison in the same way as done in Sect. III. We used a Gaussian likelihood and the same prior for $\mathcal{R}_{\mathrm{sun}}$ and $\mathcal{R}_{\mathrm{jup}}$ as used in the Bayesian regression. The null hypothesis corresponds to the ansatz that the modulation data is due to only the Sun, whereas the alternate hypothesis is that the modulation data is a combination of data from Sun and Jupiter. The Bayes factors for all the ephemerides for both epochs can be found in Table 6. The Bayes factor for 2011-2013 period is around two and hence not significant. Although the Bayes factor for the 2019-2021 period is marginally larger than with the total rate time-series data with a value of approximately four, it is still <5<5 and according to Jeffreys scale, still point to “barely worth mentioning” Trotta (2017). Therefore, a Bayesian model comparison still only provides a marginal support for an additional contribution from Jupiter to the monthly modulation data during the periods 2019-2021 and 2011-2013.